OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

remember that when a line is parallel the slope will be the same. does this help you. think about point slope form --> slope intercept form to find the equation

OpenStudy (anonymous):

no I still don't understand :( @lonnie455rich

OpenStudy (anonymous):

y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7 these are the answer choices

OpenStudy (anonymous):

I know what the answer is. I just want to help you understand. give a stab at it. write out point slope form and think about it. if you know the slope

OpenStudy (anonymous):

is it y = one thirdx + 7

OpenStudy (anonymous):

no. quit guessing. show me some work I said the slope would be the same since its parallel. perpendicular would be opposite sign reciprocal

OpenStudy (anonymous):

ooooh! so since the slope is the same the answer would be y=-3x+7

OpenStudy (anonymous):

how did you get that y intercept? show me your work

OpenStudy (anonymous):

well because I did the problem and at the end of it I went x=y/3+4 and to get rid of the y/3 I added 3+4 and got 7 then changed it back to slope interceptform y=-3x+7

OpenStudy (anonymous):

hm. what you just said is very confusing but ill put this in a form where you can see what to do and see your mistake. cause I don't understand where you got x=anything with x on the left. \[y-6=-3(x+1)\] do you see where I got this and what form it is?